Circular No. 6478

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

GX 1+4 D. Chakrabarty, Massachusetts Institute of Technology (MIT); M. H.

Finger, Universities Space Research Association; and T. A. Prince, California Institute of Technology, report for the Compton Observatory BATSE team: “The accreting x-ray pulsar GX 1+4 has been undetected by BATSE since Aug. 18, with a 90-percent-confidence upper limit of 15 mCrab (20– 60 keV) for the pulsed intensity. This is the first time that GX 1+4 has remained below our 15-mCrab detection threshold for longer than a month, since we began continuous BATSE monitoring of its pulsed flux in 1991. When last detected on Aug. 18.5 UT, the phase-averaged pulsed intensity of GX 1+4 was 29 ± 4 mCrab and the barycentric pulse frequency was (8.0381 ± 0.0001) × 10^{−3} Hz. The pulse frequency derivative changed from (−3.7 ± 0.2) × 10^{−12} Hz s^{−1} during May 28–June 12, to (0.0 ± 0.2) × 10^{−12} Hz s^{−1} during July 27–Aug. 6, suggesting that a torque reversal may have occurred in early Aug. (cf. IAUC 6105, 6153). The phase-averaged pulsed intensity flared from 59 ± 5 mCrab on June 4.5 to a maximum of 204 ± 5 mCrab on Aug. 5.5, before dropping below our detection threshold.”

W. Cui and D. Chakrabarty, MIT, write: “RXTE/PCA observations of GX 1+4 on Sept. 5.03 UT detect a phase-averaged pulsed intensity of 4.9 mCrab (2–60 keV) and a barycentric pulse frequency of (8.03156 ± 0.00042) × 10^{−3} Hz, implying that the pulsar continued to spin down after the last BATSE detection on Aug. 18, with a mean pulse frequency derivative of (−4.3 ± 0.3) × 10^{−12} Hz s^{−1}. A second RXTE/PCA observation on Sept. 25.55 did not detect any pulsations, with a 90-percent-confidence upper limit of 0.2 mCrab on the pulsed intensity. This is comparable to the low state established by EXOSAT non-detections of GX 1+4 in 1983 and 1984 (IAUC 3872). Deeper x-ray observations and optical/infrared observations of the red-giant companion are strongly encouraged.”

COMET 46P/WIRTANEN P. Lamy, Laboratoire d’Astronomie Spatiale, and his team report the

apparent detection of the nucleus of 46P with the Hubble Space Telescope (+ WFPC2) on Aug. 28: “After removing the small contribution of the (stable) coma to the central pixels, the R magnitude of the nucleus was found to vary from 21.6 to 21.9 over time intervals of 1.5 hr, though a clear rotational pattern is not evident. For a geometric albedo of 0.04, the mean effective radius is 0.58 km.”

1996 September 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green